# Causality Is Key to Understand and Balance Multiple Goals in Trustworthy ML and Foundation Models

### v4

@@ 2.2 Integrating Causality into ML
Integrating causality into ML enables models to move beyond pattern recognition and learn underlying mechanisms governing data. This section explores different approaches to causal ML, ranging from explicitly constrained models that follow predefined causal structures to methods that infer causal relationships from data.
@@
Causally Constrained ML (CCML). CCML refers to approaches that explicitly incorporate causal relationships into model training or inference as constraints or guiding principles. Given a causal graph $\mathcal{G}=(\mathbf{V},\mathbf{E})$, where $\mathbf{V}$ represents variables and $\mathbf{E}$ denotes directed edges encoding causal relationships, the goal is to ensure that the learned model $f:\mathbf{X}\to Y$ adheres to the causal structure encoded in $\mathcal{G}$ (Berrevoets et al., 2024; Zinati et al., 2024; Afonja et al., 2024; Schölkopf et al., 2016).
@@
Invariant feature learning (IFL). IFL relies on discovered implicit or latent causal features and structures. The task of Invariant Feature Learning (IFL) is to identify features of the data $\mathbf{X}$ that are predictive of the target $Y$ across a range of environments $\mathcal{E}$. From a causal perspective, the causal parents $\text{Pa}(Y)$ are always predictive of $Y$ under any interventional distribution (Kaddour et al., 2022). IFL can be achieved by regularizing the model or providing causal training data that is free of confounding.
Disentangled VAEs. VAEs aim to decompose the data $\mathbf{X}$ into disentangled latent factors $\mathcal{Z}$ that correspond to distinct underlying generative causes (Burgess et al., 2018). It can be combined with interventional experiments in mechanistic interpretability that involve “switching off” specific neurons or circuits to gain knowledge about causal workings of the complex model (Leeb et al., 2022). Causality is also used to audit models for fairness (Cornacchia et al., 2023; Byun et al., 2024) or robustness (Drenkow et al., 2024), providing insights into how decisions are influenced by sensitive variables and under distribution shifts.
@@
Double Machine Learning (DML). DML provides another causal approach by leveraging modern ML techniques for estimating high-dimensional nuisance parameters while preserving statistical guarantees in causal inference (Chernozhukov et al., 2018). DML decomposes the estimation problem into two stages: (1) predicting confounders using ML models and (2) estimating the causal effect using residualized outcomes.
@@
Causal Discovery. Finally, ML can be leveraged for causal inference or to discover causal knowledge from observational data. For instance, methods for causal discovery use statistical patterns to infer causal relationships, with notable examples including (Spirtes et al., 2000; Shimizu et al., 2006; Janzing & Schölkopf, 2010; Peters et al., 2011; Hauser & Bühlmann, 2012; Le et al., 2016).
All of the above forms a causal ML cycle (Figure 1) in which ML is enhanced by causal knowledge, controlled by causal tools, and finally contributes to enriching scientific knowledge. We include a supplementary introduction to causality and causal ML in LABEL:app:causalpreliminaries and LABEL:app:causalML.
@@ 3 Causality for Trustworthy Foundation Models
Foundation models, including state-of-the-art multimodal systems like Large Language Models (LLMs) and vision-language models, have demonstrated exceptional capabilities across diverse tasks (Achiam et al., 2023; Rohan et al., 2023; Radford et al., 2023; Brooks et al., 2024). However, their reliability remains a concern due to issues like spurious correlations, hallucinations, and unequal representation. Trade-offs and causality in trustworthy foundation models remain underexplored as compared to traditional ML. In this section, we explore the potential for causality to improve fairness, explainability, privacy, and robustness in foundation models following slightly different taxonomy than in the previous section due to their unique challenges.
@@ 3.1 Dimensions of Trustworthy Foundation Models
Fairness vs. Accuracy. Causal frameworks have become integral to fairness interventions in LLMs by identifying and mitigating pathways that lead to unfair predictions (Madhavan et al., 2023; Cotta & Maddison, 2024). Counterfactual fairness ensures that sensitive attributes (e.g., gender, race) do not causally influence outcomes. For example, in job recommendation systems, counterfactual fairness guarantees identical recommendations for equally qualified candidates regardless of their gender (Madhavan et al., 2023). Methods like causal disentanglement isolate sensitive features from output-relevant causal factors, ensuring that spurious correlations, such as gender biases in job roles, do not propagate through the model (Zhou et al., 2023; Chen et al., 2024). SCMs further enable fairness-aware fine-tuning by disentangling causal effects. However, striving for diversity has been shown to introduce non-factual output in text-to-image models. In early 2024, Google’s AI tool, Gemini, faced criticism for generating historically inaccurate images, such as depicting America’s Founding Fathers as Black individuals and Nazis as racially diverse (Vincent, 2024). Here, causality could help distinguish historically impossible scenarios from desirable diversity, ensuring both fairness and factual integrity in AI-generated content. Mode collapse is another foundation model-specific fairness issue where models generate overly generic outputs, reducing diversity and disproportionately omitting minority group representations. Causal modeling can potentially help preserve minority information by explicitly capturing causal relationships, preventing spurious correlations from erasing underrepresented patterns.
Robustness vs. Accuracy. Causal frameworks address robustness by training models to rely on invariant causal relationships while penalizing reliance on dataset-specific spurious features (Wu et al., 2024). For instance, instead of associating ”doctor” with ”male,” causal invariance enforces reliance on task-relevant features like medical terminology (Zhou et al., 2023). Causal regularization further discourages attention to non-causal patterns during inference achieving better accuracy and robustness.
@@
Privacy vs. Attribution. Causal approaches to privacy focus on detecting and severing pathways involving personally identifiable information (PII) in LLMs. Causal obfuscation uses SCMs to identify and block sensitive pathways (e.g., names, locations) during training or inference (Chu et al., 2024). Unlike traditional privacy-preserving mechanisms that indiscriminately apply noise or randomization, it ensures that only privacy-sensitive dependencies are removed, preserving essential predictive relationships.
Beyond conventional privacy concerns, attribution and memorization pose significant challenges in foundation models. Attribution is crucial in determining whether specific data—such as an artist’s work—has contributed to the training of a model, enabling rightful recognition and compensation. Memorization, on the other hand, prevents effective data removal, meaning that once a copyrighted work is embedded into a model, it becomes difficult to erase upon request. Causal auditing (Sharkey et al., 2024) potentially offers a principled way to address these challenges by providing a structured framework to verify whether a given dataset—such as an artist’s work—has influenced the model’s outputs. Unlike statistical correlation-based methods, which may falsely associate stylistic elements with broader art movements, causal auditing can disentangle direct influences from broader historical trends, ensuring that attribution is based on actual data contributions rather than incidental similarities.
Explainability vs. Capability. Although foundation models demonstrate remarkable capabilities in various tasks, their outputs often lack interpretability, making it difficult to understand or explain their reasoning. Causal models can help quantify how much each input feature contributes to a specific output, providing a clear and interpretable explanation. By modeling causal chains, we can explain how different stages of the LLM (e.g., embedding, attention layers, output logits) interact to produce a final decision (Bagheri et al., 2024). This creates a step-by-step explanation of the model’s reasoning process. Another domain that is related to causality is mechanistic interpretability. Mechanistic interpretability seeks to decode the inner workings of LLMs by analyzing their architecture, weights, and activation patterns (Conmy et al., 2023). Causality enhances this understanding by identifying cause-effect relationships within these mechanisms. Causality can identify specific pathways in neural circuits that contribute to certain outputs (Palit et al., 2023; Parekh et al., 2024). For example, specific neurons or attention heads affect token predictions, revealing the factors driving outputs.
@@ 3.2 Integrating Causality in Foundation Models
Pre-training: Causal data augmentation. Synthetic datasets with explicit causal structures, such as counterfactual examples or causal-transformable text data, can be used to augment training data. Counterfactual data augmentation introduces scenarios where causal relationships differ from spurious correlations, helping models learn true causal dependencies instead of misleading patterns (Webster et al., 2020; Chen et al., 2022).
@@
Pre-training: Causal Representation Learning. By disentangling causal factors from non-causal ones, models can learn representations that separate meaningful causal features from irrelevant associations. Techniques such as causal embedding methods (Rajendran et al., 2024; Jiang et al., 2024), which can use training data annotated with causal labels, can guide models in identifying and prioritizing true causal relationships. This has been shown to reduce reliance on spurious correlations, such as gender-biased occupational associations (Zhou et al., 2023).
@@
Pre-training: Entity interventions. SCMs can be used to intervene on specific entities (e.g., replacing “Joe Biden” with “ENTITY-A”) during pre-training (Wang et al., 2023), thus reducing entity-based spurious associations while preserving causal relationships in the data.
Pre-training: Loss function. Modifying the pre-training loss function to penalize reliance on confounders can help align models with causal principles. For instance, fine-tuning models on embeddings pre-trained with debiased token representations has shown promise for causal learning (Kaneko & Bollegala, 2021; Guo et al., 2022; He et al., 2022; Wang et al., 2023).
Post-training: Fine-tuning. Fine-tuning on datasets specifically designed to highlight causal reasoning (e.g., datasets emphasizing cause-effect linguistic patterns) ensures that models learn causal-invariant patterns. Further, counterfactual data samples can also improve the fine-tuning. Synthetic counterfactual examples improve the model’s robustness to spurious correlations, similar to pre-training, but with better sample size efficiency. Frameworks like DISCO (Chen et al., 2022) generate diverse counterfactuals during fine-tuning to enhance OOD generalization for downstream tasks. Causally Fair Language Models (CFL) (Madhavan et al., 2023) use SCM-based regularization to detoxify outputs or enforce demographically neutral generation during post-training.  Wang & Culotta (2020) use causal reasoning to separate genuine from spurious correlations by computing controlled direct effects, ensuring robust performance.
@@
Post-training: Alignment. RLHF can be adapted to include causal interventions, allowing feedback to act as instrumental variables that correct biased model behavior. Causality-Aware Alignment (CAA) (Xia et al., 2024) incorporates causal interventions to reduce demographic stereotypes during fine-tuning with alignment objectives. Extending RLHF with causal alignment to support dynamic, context-sensitive interventions could help address biases that evolve. Integrating causal reasoning into the reward model’s decision-making process, by critiquing the output of LLM using a reward model or a mixture of reward models that control for specific confounders or spurious correlations can potentially improve the downstream reasoning abilities potentially mitigating hallucinations.
Auditing and Evaluation. Causality provides a structured framework for auditing privacy risks by identifying whether sensitive user data contributes to model outputs. This is particularly important for privacy regulations like GDPR’s “right to be forgotten” (EC, 2016), where users can request their data to be removed from an AI system. However, verifying whether an LLM has truly forgotten a user’s data is a complex challenge, as models can memorize training information in ways that are difficult to detect through standard evaluation metrics. One key approach in privacy auditing is using causal attribution, which assesses whether a specific data point influenced a given output. By using do-calculus, privacy auditors can evaluate how an output changes when a particular data source is removed. This enables a principled test of whether an LLM has truly forgotten a user’s data.
Practical Considerations. In supervised fine-tuning and alignment, the downstream task and its causal relationships are often known, allowing for more targeted interventions on confounding variables and even the collection of task-specific data to refine causal structures. Additionally, since post-training typically requires less data than pre-training, integrating causal insights becomes more feasible. Pre-training offers the advantage of learning broad representations from diverse data, but it is difficult to enforce causal constraints due to the lack of explicit task definitions and causal structures. Auditing is particularly useful for detecting biases, ensuring fairness, and validating robustness in real-world scenarios. Unlike pre-training and fine-tuning, auditing does not require modifying the training pipeline, making it a cost-effective way to introduce causal reasoning retrospectively.
@@ 4 Challenges and Opportunities
Despite its advantages, there are many challenges when applying causality to trustworthy ML, including reliance on strong causal assumptions and limited availability of a priori causal knowledge, particularly in the form of DAGs. Foundation models bring further complications due to their scale, high-dimensional data, and the difficulty of validating causal structures. We outline key obstacles in integrating causality into ML and foundation models and suggest strategies to overcome them.
@@
Availability of Causal Knowledge. A major challenge in causal ML is the limited availability of causal knowledge, particularly in the form of DAGs. Expert-constructed DAGs may suffer from subjectivity and scalability issues, while ML-based causal discovery is constrained by identifiability assumptions and noise sensitivity. However, recent hybrid approaches combining classical causal discovery with LLM-based reasoning offer promising solutions.
@@
Causal Transportability. Scientific knowledge often lacks direct applicability across different populations, making causal transportability essential. Pearl and Bareinboim’s DAG-based framework adjusts causal knowledge for new settings using targeted data collection (Pearl & Bareinboim, 2011b; Bareinboim & Pearl, 2014; Pearl & Bareinboim, 2011a). Building on this, Binkyte et al. (2024) propose an expectation-maximization (EM) approach to adapt causal knowledge for target demographic applications.
Potentially Unresolvable Tensions. Not all tensions in trustworthy AI can always be fully resolved. For instance, stronger privacy protections often reduce model utility (Dwork et al., 2014; Bassily et al., 2014). Similarly, explainability may sometimes come at the cost of accuracy, and robustness can conflict with fairness in certain scenarios. However, causality provides a structured approach to evaluating these trade-offs, making it possible to quantify their impact and identify cases where full reconciliation is not feasible. Importantly, it is crucial to be transparent about these limitations, as this fosters societal trust, promotes accountability, and enables more informed decision-making in AI development.

### v5

@@ 2.2 Integrating Causality into ML
Integrating causality into ML allows models to move beyond surface-level patterns and capture the underlying mechanisms that generate data. This section introduces methods for incorporating causal knowledge into ML models, as well as techniques for using ML to discover causal relationships from data.
@@
Causally Constrained ML (CCML). CCML refers to approaches that explicitly incorporate causal relationships into model training or inference as constraints or guiding principles. Given a causal graph $\mathcal{G}=(\mathbf{V},\mathbf{E})$, where $\mathbf{V}$ represents variables and $\mathbf{E}$ denotes directed edges encoding causal relationships, the goal is to ensure that the learned model $f:\mathbf{X}\to Y$ adheres to the causal structure encoded in $\mathcal{G}$ berrevoets2024causal ; zinati2024groundgan ; afonja2024llm4grn ; Scholkopfetal16 .
@@
Invariant feature learning (IFL). IFL relies on discovered implicit or latent causal features and structures. The task of Invariant Feature Learning (IFL) is to identify features of the data $\mathbf{X}$ that are predictive of the target $Y$ across a range of environments $\mathcal{E}$. From a causal perspective, the causal parents $\text{Pa}(Y)$ are always predictive of $Y$ under any interventional distribution kaddour2022causal . IFL can be achieved by regularizing the model or providing causal training data that is free of confounding.
Disentangled VAEs. VAEs aim to decompose the data $\mathbf{X}$ into disentangled latent factors $\mathcal{Z}$ that correspond to distinct underlying generative causes, especially when trained with causal constraints or interventional data burgess2018understanding ; kim2021counterfactual . This aligns with causal reasoning, as each latent dimension can be interpreted as an independent causal factor influencing the observed data.
@@
Double Machine Learning (DML). DML provides another causal approach by leveraging modern ML techniques for estimating high-dimensional nuisance parameters while preserving statistical guarantees in causal inference chernozhukov2018double . DML decomposes the estimation problem into two stages: (1) predicting confounders using ML models and (2) estimating the causal effect using residualized outcomes.
@@
Causal Discovery. Finally, ML can be leveraged for causal inference or to discover causal knowledge from observational data. For instance, methods for causal discovery use conditional independence, score, or noise asymmetry-based methods, to infer causal relationships, with notable examples including spirtes2000causation ; shimizu2006linear ; janzing2010causal ; PetMooJanSch11 ; hauser2012gies ; le2016fast . More recent methods leverage Deep Neural Networks (DNNs) in combination for causal discovery nauta2019causal ; bi2023large ; goudet2017causal .
Summary and Use Cases.
The methods presented above represent a spectrum of causal integration strategies in ML. CCML requires explicit or semi-explicit causal structures, offering high interpretability but often demanding strong prior knowledge. However, explicit causal guardrails are beneficial for high-stakes applications requiring clear accountability. In contrast, IFL, DML, and Disentangled VAEs operate with weaker structural assumptions and are well-suited for settings with limited causal annotations as well as exploratory tasks or scalable representation learning in unstructured data. Causal Discovery lies at the other end of the spectrum, aiming to uncover causal relationships from data alone. Combining methods—for example, using causal discovery to inform CCML also help balance rigor and practicality.
@@ 3 Causality for Trustworthy Foundation Models
Foundation models (FMs), including state-of-the-art multimodal systems like LLMs and vision-language models, have demonstrated exceptional capabilities across diverse tasks achiam2023gpt ; team2023gemini ; radford2023robust . However, their reliability remains a concern due to issues like spurious correlations, hallucinations, and unequal representation. Trade-offs and causality in trustworthy foundation models remain underexplored as compared to traditional ML. In this section, we explore the potential for causality to improve fairness, explainability, privacy, and robustness in foundation models following a slightly different taxonomy than in the previous section due to their unique challenges.
@@ 3.1 Dimensions of Trustworthy Foundation Models
Fairness vs. Accuracy. Causal frameworks support fairness in LLMs by identifying and mitigating pathways that lead to biased predictions d011b9cc8bad2d29595f1ef5cff344abd2c5c5ab ; c9db178549df34ea26994753e3560afac761f39e . Techniques like counterfactual fairness ensure that sensitive attributes (e.g., gender, race) do not influence outcomes, while causal disentanglement separates spurious correlations from true predictive factors zhou2023causal ; 3a2e333502f5035f0f78318c7722ec1a7b7265bb . Structural causal models (SCMs) further enable fairness-aware fine-tuning by isolating causal effects d011b9cc8bad2d29595f1ef5cff344abd2c5c5ab .
In addition, in generative models, careless fairness interventions can backfire—for instance, efforts to enhance diversity have led to historically inaccurate outputs, as seen in critiques of Google’s Gemini verge2024gemini . Causal reasoning can help balance fairness and factuality by distinguishing meaningful diversity from implausible scenarios. It also offers tools to address mode collapse, preserving minority representation by capturing underrepresented causal patterns.
@@
Privacy vs. Attribution. Causal approaches to privacy focus on identifying and blocking pathways involving personally identifiable information (PII) in LLMs. Causal obfuscation uses SCMs to remove only privacy-sensitive dependencies (e.g., names, locations), preserving essential predictive signals chu2024causal .
Attribution and memorization raise further concerns. Attribution determines whether specific data, such as an artist’s work, contributed to model training, while memorization complicates data removal. Causal auditing sharkey2024causal offers a principled alternative to correlation-based methods by disentangling direct data influences from incidental stylistic similarities, enabling more accurate and accountable attribution.
Robustness vs. Accuracy. Causal frameworks address robustness by training models to rely on invariant causal relationships while penalizing reliance on dataset-specific spurious features 9f597bdbd47d38e3920d2ab9020b91a558cdf026 . For instance, instead of associating "doctor" with "male," causal invariance enforces reliance on task-relevant features like medical terminology zhou2023causal . Causal regularization further discourages attention to non-causal patterns during inference, achieving better accuracy and robustness.
Explainability vs. Capability. Despite their impressive capabilities, foundation models often lack interpretability, making their reasoning difficult to explain. Causal models address this by quantifying how input features influence outputs and modeling interactions across the model’s components (e.g., embeddings, attention, logits)bagheri2024c2p . This enables step-by-step explanations of the model’s decisions. Closely related is mechanistic interpretability, which analyzes model architecture, weights, and activations conmy2023towards . Interventional methods enhance this by uncovering cause-effect relationships within neural circuits palit2023towards ; parekh2024concept . Application of formal causal theory and methods to mechanistic interpretability is an emerging research direction that shows promise for deeper, more robust insights geiger2024causal ; stolfo2023mechanistic ; sharkey2025open ; sharkey2024causal .
@@ 3.2 Integrating Causality in Foundation Models
Causal Representation Learning. By disentangling causal factors from non-causal ones, models can learn representations that separate meaningful causal features from irrelevant associations. Techniques such as causal embedding methods rajendran2024learning ; jiang2024origins , use training data annotated with causal labels. This has been shown to reduce reliance on spurious correlations, such as gender-biased occupational associations zhou2023causal . Fine-tuning models on embeddings pre-trained with debiased token representations has shown promise for causal learning kaneko2021debiasing ; guo2022auto ; he2022mabel ; wang2023causal .
@@
Entity interventions. SCMs can be used to determine interventions on specific entities (e.g., replacing “Joe Biden" with “ENTITY-A") during pre-training wang2023causal , thus reducing entity-based spurious associations while preserving causal relationships in the data.
Counterfactual Data Augmentation. Synthetic datasets with explicit causal structures or counterfactual examples, that introduce scenarios where causal relationships differ from spurious correlations, help models learn true causal dependencies instead of misleading patterns webster2020measuring ; chen2022disco . Synthetic causal data can also be used for fine-tuning the models, similar to pre-training, but with better sample size efficiency. Frameworks like DISCO chen2022disco generate diverse counterfactuals during fine-tuning to enhance OOD generalization for downstream tasks.
SCM-based Methods. Causally Fair Language Models (CFL) d011b9cc8bad2d29595f1ef5cff344abd2c5c5ab use SCM-based regularization to detoxify outputs or enforce demographically neutral generation during post-training.  wang2020identifying use causal reasoning to separate genuine from spurious correlations by computing controlled direct effects, ensuring robust performance.
@@
Causal RLHF Alignement. RLHF can be adapted to include causal interventions, allowing feedback to act as instrumental variables that correct biased model behavior. Causality-Aware Alignment xia-etal-2024-aligning incorporates causal interventions to reduce demographic stereotypes during fine-tuning with alignment objectives. Extending RLHF with causal alignment to support dynamic, context-sensitive interventions could help address biases that evolve. Integrating causal reasoning into the reward model’s decision-making process, by critiquing the output of LLM using a reward model or a mixture of reward models that control for specific confounders or spurious correlations, can potentially improve the downstream reasoning abilities, potentially mitigating hallucinations.
Latent Representation Discovery. Recent work in mechanistic interpretability has employed sparse autoencoders to extract interpretable latent representations aligned with internal circuit features kissane2024interpreting ; conmy2023towards . These methods enable the identification of functional substructures within models by isolating meaningful directions in activation space.
Targeted Interrogation Methods. Techniques based on interventions, such as ablation, activation patching, or causal scrubbing, allow us to move beyond passive observation by directly testing the functional roles of model components brinkmann2024mechanistic ; syed2023attribution . These methods provide a powerful framework for identifying how specific neurons or pathways contribute to behavior, enabling rigorous, theory-driven interpretability.
Causal Discovery. The recent advancements in large language models (LLMs) have inspired their use in causal discovery (kiciman2023causal, ; kasetty2024evaluating, ; vashishtha2023causal, ; ai4science2023impact, ; abdulaal2023causal, ; khatibi2024alcm, ). Most of the above methods involve the refinement of the statistically inferred causal graph by LLM. However, emerging research shows that LLMs can be useful in constructing full causal graphs based on parametric knowledge of scientific literature in diverse domains (sheth2024hypothesizing, ; afonja2024llm4grn, ).
Summary and Use Cases.
These methods span both causal inference and integration. Causal data augmentation, representation learning, and entity interventions are methods used in pretraining, where broad structural priors can be encoded. In contrast, causal fine-tuning, alignment, and auditing are post-training methods requiring less causally annotated data, but more explicit goals and causal structures. Latent representation discovery and targeted interrogation methods stand out as a retrospective techniques that can help to assess fairness, privacy, and robustness without modifying the training pipeline. Causal discovery can support both pre- and post-training by informing model design or evaluating learned representations. Selecting among these approaches depends on the availability of causal knowledge, the interpretability needs, and the constraints of the training pipeline.
@@ 4 Challenges and Opportunities
Despite the advantages of a causal approach, practical applications pose challenges, including reliance on strong causal assumptions and limited availability of a priori causal knowledge. FMs bring further complications due to their complexity and scale. We outline key obstacles in integrating causality into ML and FMs and suggest strategies to overcome them.
4.1 General Challenges in Applying Causality
@@
Availability of Causal Knowledge. A major challenge in causal ML is the limited availability of causal knowledge, particularly in the form of DAGs. Expert-constructed DAGs may suffer from subjectivity and scalability issues, while ML-based causal discovery is constrained by identifiability assumptions and noise sensitivity. However, recent hybrid approaches combining classical causal discovery with LLM-based reasoning offer promising solutions. In addition, approximate causal interventions are possible without fully specified model or with partial causal graph russell2017worlds ; ehyaeibridging ; zuo2022counterfactual .
@@
Causal Transportability. Scientific knowledge often lacks direct applicability across different populations, making causal transportability essential. Pearl and Bareinboim’s DAG-based framework adjusts causal knowledge for new settings using targeted data collection bareinboim2014transportability ; pearl2011transportability . Building on this, binkyte2024babe propose an expectation-maximization (EM) approach to adapt causal knowledge for target demographic applications.
Potentially Unresolvable Tensions. Not all tensions in trustworthy AI can always be fully resolved. For instance, stronger privacy protections often reduce model utility dwork2014algorithmic ; bassily2014private . We also acknowledge that some fairness conflicts stem from deeper normative or value-based disagreements, for example, a causal relationship may exist, but relying on it in decision-making may still be viewed as unfair from an ethical or legal standpoint. In such cases, causality does not offer a definitive solution but provides a framework to surface these assumptions transparently, enabling more informed discussions that go beyond technical considerations.
4.2 Challenges in Causal Foundation Models